OpenStudy (anonymous):

The sum of two numbers is 86. One of the numbers is five less than three times the other number. What are the two numbers?

OpenStudy (anonymous):

x + y = 86 x = 3y - 5 3y - 5 = 86 - y

OpenStudy (anonymous):

Solve for y

OpenStudy (anonymous):

Would you add 5 to both sides first?

OpenStudy (anonymous):

3y-5=86-y =>4y=91 =>y=91/4

OpenStudy (anonymous):

Thanks Rahul.....

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

I was asking lilangl to solve for y

OpenStudy (anonymous):

sry...then

OpenStudy (anonymous):

I dont understand?

OpenStudy (anonymous):

y = 22.75

OpenStudy (anonymous):

if you solve for y, that's what it comes out to

OpenStudy (anonymous):

You were on the right track lilangl

OpenStudy (anonymous):

would i have to convert it to a decimal to determine the other number?

OpenStudy (anonymous):

sorry, didnt see that you already did that...lol

OpenStudy (anonymous):

so then you would take the second equation of x = 3y - 5 and substitute the 22.75 for y and solve for x?

OpenStudy (anonymous):

\[3y - 5 = 86 - y \]\[3y - 5 + 5= 86+5 - y\]\[3y = 91 - y\]\[3y + y= 91 - y +y\]\[4y = 91\]\[4y/4 = 91/4\]\[y = 91/4\]\[y = 22.75\]

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

Yes, but you could also insert it into the first equation.

OpenStudy (anonymous):

x + 22.75 = 86

OpenStudy (anonymous):

And then solve for x that way

OpenStudy (anonymous):

:o)

OpenStudy (anonymous):

What is x?

OpenStudy (anonymous):

67.25

OpenStudy (anonymous):

Hmmm, try again....

OpenStudy (anonymous):

You were close